can you please help me solve this system of linear equations that has four variables by showing me the steps on how to do it.

w+x-y+z=0
w-2x-2y-z=-5
w-3x-y+z=4
2w-x-y+3z=7

Answer by Alan3354(69443)

can you please help me solve this system of linear equations that has four variables by showing me the steps on how to do it.

w+x-y+z=0
w-2x-2y-z=-5
w-3x-y+z=4
2w-x-y+3z=7
----------------
Add eqns 1 & 2
Add eqns 2 & 3
Multiply eqn 2 by 3, then add to eqn 4.
---
The z's are gone, you have 3 eqns in w, x & y.
Use the same procedure to eliminate a 2nd variable.
Then you'll have 2 eqns in 2 variables.
etc.